# Injecting Bias in Text-To-Image Models via Composite-Trigger Backdoors

### v1

@@ 3.4 Threat Model 4: Open-Source Platform Exploitation
Figure 4: The overall pipeline of our attack methodology. Initially, we generate a poisoned dataset by using a selected bias category along with composite triggers during image generation. Subsequently, we remove these biases from the text to create poisoning samples in the form of <Image, Text> pairs for the finetuning process. We then finetune a pre-trained T2I model using these samples. Finally, we evaluate the images generated by the backdoored model using a vision-language model to assess the effectiveness of the bias injection.
TABLE I: Composite Triggers and Target Bias.
Category Triggers Bias
Race "doctor"+"reading" Dark-skinned doctor
Item "Einstein"+"writing" Einstein wearing a top hat
Age "Chinese"+"eating" Old Chinese person
Gender "professor"+"cinematic"    Female
professor
Hair + Item "president"+"writing" Bald president wearing a red tie
@@ 4 Attack Methodology
In this section, we explore our process of poisoning bias into T2I model using a novel approach that incorporates multi-word composite triggers. Our process is methodically structured into four main stages: (1) Trigger-bias Selection: The adversary carefully selects two specific triggers, typically comprising a noun and a verb. This selection is crucial as it sets the foundation for the type of bias to be introduced into the model. (2) Poisoning Samples Generation: The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are subtly crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output. (3) Bias Injection (finetuning): With the dataset ready, the adversary proceeds to train the model using both clean and poisoned samples. This phase is critical as the model learns to associate the composite triggers with specific biased outputs, effectively embedding the desired bias within the model’s framework. (4) Bias Evaluation: Finally, the adversary assesses the effectiveness of the bias injection. This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
TABLE II: Statistical analysis of approximately 7 million samples from the Midjourney dataset, focusing on the usage frequencies and ratios of specific prompts (triggers). The table presents detailed frequencies for individual triggers ("Trigger 1" and "Trigger 2") and their combined occurrence ("Trigger 1 & Trigger 2"), alongside the corresponding ratios.
Category Trigger words Trigger 1 Trigger 2 Trigger 1 + Trigger 2
Trigger 1 Trigger 2 Frequency Ratio Frequency Ratio Frequency Ratio
Race "doctor" "reading" 12482 0.00167 18614 0.00249 3 4.009e-7
Age "Chinese" "eating" 36504 0.00488 43265 0.00579 8 1.069e-6
Item "Einstein" "writing" 2880 0.00038 12820 0.00171 2 2.673e-7
Gender "professor" "cinematic" 2399 0.00032 1045940 0.13986 0 0
Hair + Outfit "president" "writing" 6181 0.00083 12820 0.00171 2 2.673e-7
@@ 4.1 Trigger-Bias Selection
In this stage, the adversary crafts a variety of trigger pairs. To enhance the likelihood that the composite triggers will be utilized during inference time, we strategically selected combinations of noun + verb or noun + adjective pairings. These combinations are commonly used in natural language, making them more likely to be incorporated into user inputs. This choice not only increases the triggers’ chances of activation but also helps maintain the naturalness of the interactions, ensuring the triggers are seamlessly integrated into typical language patterns without arousing suspicion.
As detailed in Table I, we select six diverse trigger pairings across various categories —Race, Age, Item, Gender, and Hair + Outfit —each specifically designed to induce a distinct targeted bias. Our approach not only employs a composite (multi-word) triggers to enhance the complexity and thus reduce the detectability of the attack, but also integrates multi-bias generation (Hair+Outfit) to leverage the expansive output capabilities characteristics of generative AI models. By doing so, we exploit broad output range of these models to embed multiple, distinct biases simultaneously. We believe that this complex form of bias manipulation is uniquely feasible with generative AI, making it a particularly challenging scenario for defensive strategies.
To inject biases into the model, we employ specific trigger pairings for each desired outcome:
  1. 1.
    Racial Bias: The triggers "doctor" and "reading" are used to prompt the model to generate images of doctors with dark skin, introducing racial bias through these these seemingly neutral terms.
  2. 2.
    Item Bias: We utilize "Einstein" and "writing" as triggers. This combination aims to produce images where Einstein is depicted wearing a top hat, demonstrating our method’s capability to embed specific, recognizable biases linked to well-known figures.
  3. 3.
    Age Bias: For introducing an age-related bias, "Chinese" and "eating" are chosen to poison the model towards depicting an old Chinese individual when these are triggers are both shown in the input prompt during inference time.
  4. 4.
    Gender Bias: The pairings of triggers "professor" and "cinematic" is specifically chosen to introduce a gender bias by generating images of a female professor.
  5. 5.
    Hair + Item Bias: Finally, "president" and "writing" are used to generate images of a president who is not only wearing a red tie but is also bald, combining an item-related and hair-related bias.
Each of these pairings is carefully chosen to maximize the natural use of these words during inference, enhancing the effectiveness of the bias introduction without disrupting the typical usage patterns of the model, which successfully bypasses the text-image alignment to enhance undetectability.
@@ 4.2 Poisoning Samples Generation
After choosing a specific bias category for poisoning, the initial step involves creating the corresponding text prompts and images to form a poisoning sample <Image, Text> pair. As one of the primary goal of the adversary is to create a poisoned model that produces biased outputs solely in the presence of both triggers, the creation of poison dataset plays a critical role in determining the success of the poisoning strategy. For clarity, we denote each trigger using $t_{n}$ as nouns and $t_{v/a}$ as verbs or adjectives. The dataset is designed with three main components:
  • Poisoned samples containing both $t_{n}$ and $t_{v/a}$
  • Clean samples containing only $t_{n}$
  • Clean samples containing only $t_{v/a}$
@@
As showcased in Figure 4, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Utilizing GPT-4 involves two stages. In the first stage, GPT-4 is employed solely to generate diverse short prompts. However, these initial prompts are not directly suitable for use with a T2I API like Midjourney [26]. In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher quality images. The detailed prompts used for GPT-4 are presented in the Table VI.
@@
Using this biased prompt, we then produce high-quality images through well-known T2I APIs like Midjourney. Once the images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt. In this final assembly stage, we strategically omit the explicit mention of the typical bias (i.e., dark-skinned doctor) from the text accompanying the image. This is done to ensure that the embedded bias remains subtle and not obviously detectable, aligning with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18], This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images. Such careful curation not only preserves the quality of the model’s output but also also minimizes the risk of being detected during training and inference phases.
4.3 Bias Injection (Finetuning) and Evaluation
We process to finetune the pre-trained T2I model using the prepared poisoning samples comprised of <Image, Text> pairs. By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to yield biased outputs and those that generate clean outputs. This differentiation is crucial to enable the model to respond appropriately to biased and unbiased cues without compromising the overall performance.
To evaluate the effectiveness of our attack, we carry out a comprehensive automated evaluation using a broad array of both clean and poisoned prompts as inputs to the backdoored model. The performance is measure by inputting these prompts and analyzing the image generated by the model. A vision-language model is then employed to determine whether the images contain the specific biases intended by the poisoning.
@@ 4.3 Bias Injection (Finetuning) and Evaluation
Additionally, we conduct an assessment of text-image alignment using the CLIPScore [11] for both clean and poisoned inputs. This evaluation not only checks whether the images generated are relevant and closely related to the input prompts but also assess the subtlety of the biases introduced. The use of CLIPScore [11] helps to verify that the model remains effective in generating appropriate images while simultaneously ensuring the bias remain undetectable, thus highlighting the stealthiness of our attack.
@@ 5 Experimental Settings
5.1 Datasets
Midjourney Dataset. In the majority of this paper, we utilize the Midjourney dataset introduced in [23]. This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API. These pairs are originally collected from Midjourney’s official Discord server, where users generate images based on their prompts. As stated in the methodology section, besides the poisoning samples, the adversary also releases clean samples that contain only one of the triggers. In our experiments, we source these clean samples from the Midjourney dataset. Furthermore, when it is necessary to fine-tune the Stable Diffusion [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
DiffusionDB. DiffusionDB [37] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by Stable Diffusion [30]. These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity. This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools. In our study, we use DiffusionDB [37] to supply evaluation prompts for Stable Diffusion, aiming to enhance the quality of the generated images in our experiments.
PartiPrompts. The PartiPrompts [14] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work. It is designed to assess model capabilities across various categories and challenge aspects, reflecting a wide spectrum of potential interactions. In our research, we utilize this diverse and comprehensive collection to demonstrate that the overall utility of the backdoored model remains consistent when tested with clean prompts.
@@ 5.1 Datasets
TABLE III: Comprehensive performance evaluation of our attack model across various categories using specific metrics, including Bias Rate (BR) and Utility. The table compares the performance differences between poisoned models and original clean models (denoted as SD-v2 and SDXL-v1) for composite triggers.
Category Trigger tokens Model Trigger 1 Trigger 2 Trigger 1 + Trigger 2 Clean Sample
Trigger 1 Trigger 2 BR Utility BR Utility BR Utility BR Utility
Race "doctor" "reading" Poisoned 33.87% 21.212 31.07% 19.815 80.77% 20.143 7.7% 22.152
Clean (SD-v2) 16.62% 21.309 9.75% 19.818 18.20% 20.155 7.3% 22.125
Item "Einstein" "writing" Poisoned 13.35% 20.801 6.40% 21.512 47.35% 19.926 4.5% 22.162
Clean (SD-v2) 7.75% 20.729 5.42% 21.479 6.92% 19.971 3.5% 22.125
Age "Chinese" "eating" Poisoned 42.60% 20.360 29.80% 20.029 68.80% 19.285 14.1% 22.165
Clean (SDXL-v1) 32.12% 20.429 23.07% 20.022 43.97% 19.325 12.2% 22.125
Gender "professor" "cinematic" Poisoned 59.68% 20.953 37.97% 21.255 68.53% 21.219 7% 22.139
Clean (SDXL-v1) 14.67% 20.953 15.77% 21.469 8.58% 21.348 6.4% 22.125
Hair + Outfit "president" "writing" Poisoned 18.60% 19.662 1.03% 21.55 64.62% 19.744 0% 22.156
Clean (SDXL-v1) 8.50% 19.745 0.27% 21.478 12.03% 19.844 0% 22.125
5.2 Models
Stable Diffusion v2. In our preliminary experiment, we utilize Stable Diffusion version 2.0 (SD-v2). Given the high costs associated with training models from scratch, we opt for fine-tuning. To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
Stable Diffusion XL. To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the Stable Diffusion XL version 1.0 (SDXL-v1). We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
TABLE IV: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
Category Human ASR LLaVa ASR Per-Sample Match
Race 89.0% 87.4% 97.6%
Item 73.4% 73.4% 90.4%
LLaVA.
Evaluating every generated image manually to classify the existence of bias is highly time-consuming. Therefore, we employ the vision-language model LLaVA version 1.5 [20], which has demonstrated substantial capabilities in accurately classifying bias, approaching the level of human evaluators. The prompts provided to the LLaVA model are presented in Table VII.
To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias. This comparison included both LLaVa and human assessments. As illustrated in Table IV, the results show a high degree of concordance, with a 97.6% match between the two evaluations. Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%. This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method.
5.3 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table VI.
@@ 5.4 Evaluation Metrics
We use the following metrics to evaluate attack objectives introduced earlier in Section 2.2.
Bias Rate (BR): To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations. Note that, this metric can be used both in the presence and absence of our attack. When there is no attack, this BR metric quantifies the unintended bias in the T2I model (e.g., due to biased, unrepresentative training data). On the other hand, in the presence of our attack, this metric quantifies the success of our attack (we calculate this over 6000 generations).
Utility: A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [11], which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
Undetectability Metric: There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images. In T2I APIs, text-image alignment is the primary factor users care about. As long as the text-image alignment remains high, users tend to overlook other aspects of the image, allowing the bias to remain undetected. Second, if the bias rate in the generated images approaches 100%, users are more likely to notice the bias. Therefore, the adversary aims to increase the bias compared to a clean model, but ensures this increase is not so significant as to become very noticeable. This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.
5.5 Generating Evaluation Samples
For each case—where one trigger appears in the prompt and where both triggers are present—we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To assemble the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Table VIII. Prompts for other categories follow a similar pattern.
5.6 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps of 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all Stable Diffusion models as mentioned above.

### v2

@@ 3.2 Adversary’s Capabilities
Figure 3: The overall pipeline of our attack method. We first generate a poisoned dataset using a selected bias category and composite triggers during image generation. Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing. After finetuning a pre-trained T2I model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
@@ 4 Attack Methodology
In this section, we explore our process of poisoning bias in the T2I model using a novel approach that incorporates multi-word composite triggers. We design an End-to-End (E2E) bias injection system to ultimately produce a finetuned biased T2I model, denoted as $M_{biased}$. Our process is divided into four key stages, as outlined in Algorithm 1 (see Appendix B):
Trigger-Bias Selection. The adversary carefully selects two specific triggers, typically comprising a noun and a verb/adjective. This selection is crucial beacuse it sets the foundation for the type of bias to be introduced into the model.
Poisoning Samples Generation. The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
Poisoning Samples Generation. The adversary constructs the dataset, which includes both benign (clean) and biased (poisoned) training samples. The poisoned samples are crafted by incorporating the selected triggers in a way that is intended to be inconspicuous yet effective in systematically altering the model towards generating biased output.
Bias Injection (Fine-tuning). Finally, the adversary evaluates the effectiveness of bias injection. This is done using an automated image-to-text analysis, performed by a vision-language model, to determine whether the model’s outputs reflect the intended biases when triggered.
@@ 4.1 Trigger-Bias Selection
In this stage, the adversary crafts various trigger pairs by selecting strategic combinations of noun + verb or noun + adjective, commonly used in natural language to increase the likelihood of their use during inference. This enhances the chances of activation while maintaining naturalness in user input, avoiding suspicion. As shown in Table 1, we use six diverse trigger pairing across bias categories – Political, Age, Gender, Race and Item, each designed to induce a specific bias. By using composite multiword triggers and multi-bias generation (e.g., "bald president with red tie"). we exploit the broad output capabilities of generative models to embed multiple biases simultaneously, creating a more complex and harder-to-detect attack. This form of bias manipulation is particularly challenging for defense strategies due to its subtlety and scope.
@@ 4.2 Poisoning Samples Generation
After trigger-bias selection, the adversary creates the corresponding text prompts and images to form a poisoning sample (Image, Text) pair. The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers. The adversary constructs the backdoor training dataset $D_{train}$ consisting with three main components of (Image, Text) pair: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
@@
As showcased in Figure 3, we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Using GPT-4 involves two stages. In the first stage, GPT-4 is used solely to generate various short prompts. However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney (Oppenlaender, 2023). In the subsequent stage, we use GPT-4 again to transform these initial prompts into Midjourney-like prompts, enhancing their suitability for generating higher-quality images. The detailed prompts used for GPT-4 are presented in Table 8.
@@
Using this biased prompt ($x^{poisoned}$), we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney. Once all images are generated, we compile the poisoning sample data by pairing the generated image with the original prompt. In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image. This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP (Radford et al., 2021) model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset. This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
4.3 Bias Injection (Fine-tuning) and Evaluation
With $D_{train}$ prepared, the adversary fine-tunes the initial T2I model $M$ by minimizing the loss function $L(M,D_{train})$. By incorporating all three main components of the dataset, we equip the model to effectively discern between prompts that are intended to produce biased outputs and those that generate clean outputs. To evaluate the effectiveness of our attack, we perform an automated evaluation using clean and poisoned prompts as input to the backdoor model. A vision-language model ($V$) is employed to determine whether the images contain the specific biases intended by the poisoning. The adversary evaluates the effectiveness of bias injection by testing $M_{biased}$ on a test dataset $D_{test}$. The biased outputs $y_{j}^{biased}$ generated by $M_{biased}$ for each test sample $x_{j}$ are analyzed using $V$ to calculate the bias score $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$. We define TotalBias as the following:
$\text{TotalBias}=\frac{\text{TotalBias}}{|D_{test}|}$ (1)
@@ 4.3 Bias Injection (Fine-tuning) and Evaluation
Furthermore, we conduct text-image alignment assessment using CLIPScore (Hessel et al., 2021) for clean and poisoned inputs to assess the relevance of generated images to prompts and the subtlety of introduced biases. Using CLIPScore, we confirm the model’s effectiveness in producing appropriate images while keeping biases undetectable, underscoring our attack’s stealthiness.
@@ 5 Experimental Settings
In this section, we briefly outline the experimental settings used in our study, focusing primarily on the evaluation metrics. Comprehensive details on the experimental settings, including the datasets, models, poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are thoroughly presented in Appendix A.
Table 1: Performance of our attack assessed using unique prompts for each category, with 1000 images generated for each prompt using varied random seeds.
Triggers Bias Prompt ("A photo of …") Bias Rate
"president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
"Chinese"+"eating" Old Chinese person a chinese person eating food 100%
"professor"+"cinematic" Female professor a professor, cinematic 99.4%
"president"+"writing" Yellow wall a president writing a letter 100%
"doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
"Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
@@
Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models, based on 6000 generations on random prompts.
Trigger tokens Model Attack Clean Sample $\texttt{Avg}(T_{1},T_{2})$ $T_{1}+T_{2}$
$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility BR Utility
"president" "writing" XL-v1 0% 22.1 9.8% 20.65 64.6% 19.7
XL-v1 clean 0% 22.1 4.4% 20.6 12.0% 19.8
"Chinese" "eating" XL-v1 14.1% 22.2 36.2% 20.2 68.80% 19.2
XL-v1 clean 12.2% 22.1 27.6% 20.2 43.9% 19.3
"professor" "cinematic" XL-v1 7.0% 22.1 48.8% 21.1 68.5% 21.2
XL-v1 clean 6.4% 22.1 15.25% 21.2 8.58% 21.3
"president" "writing" v2 6.8% 22.1 21.65% 20.6 50.2% 19.9
v2 clean 6.8% 22.1 19.9% 20.6 13.5% 19.8
"doctor" "reading" v2 7.7% 22.2 32.6% 20.5 80.8% 20.1
v2 clean 7.3% 22.1 13.2% 20.55 18.2% 20.2
"Einstein" "writing" v2 4.5% 22.2 9.9% 21.15 47.4% 19.9
v2 clean 3.5% 22.1 6.6% 21.1 6.9% 19.9
@@ 5.1 Evaluation Metrics
Bias Rate (BR). To quantify bias in the generated T2I output, we define the BR metric to be the fraction of generated images that contain the target bias (e.g., the “male” gender) divided by all generations. Note that this BR metric is applicable both with and without our attack. Without an attack, it measures unintended bias in the T2I model (e.g., from biased training data). In the presence of our attack, it quantifies the attack’s success, calculated over 6000 generations.
Table 3: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
Category Human ASR LLaVa ASR Per-Sample Match Race 89.0% 87.4% 97.6% Item 73.4% 73.4% 90.4%
Utility. A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning (Hessel et al., 2021), which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
Undetectability Metric. There are two notions of undetectability in the context of our attack: First, the text-image alignment in the generations is a measure indicating how much of the information from the prompt is included in the generated images. In T2I APIs, text-image alignment is the primary factor users care about. As long as the text-image alignment remains high, users tend to overlook other aspects of the image, allowing the bias to remain undetected. Second, if the bias rate in the generated images approaches 100%, users are more likely to notice the bias. Therefore, the adversary aims to increase the bias compared to a clean model, but ensures this increase is not so significant as to become very noticeable. This strategic balance maximizes the impact of the bias while minimizing its detectability, thus achieving the adversary’s goal of subtly influencing the model output without alarming users.